the width of a rectangle is 15 ft. what is the perimeter of the rectangle if the length is 5 ft longer than the width?
Answer by drk(1908)

w = 15
L = w + 5 = 20
P = 2W + 2L
P = 2*15 + 2*20 = 30 + 40 = 70 feet.